OpenStudy (anonymous):

The altitude of a triangle is increasing at a rate of 2.0 centimeters/minute while the area of the triangle is increasing at a rate of 4.0 square centimeters/minute. At what rate is the base of the triangle changing when the altitude is 10.5 centimeters and the area is 84.0 square centimeters?

OpenStudy (anonymous):

i guess this was not clear the first time, so lets go slow. you know \[A'(t)=40, h'(t)=2\] you also know \[A=\frac{1}{2}bh\] that is Area is one half base times height. you want \[b'(t)\]

OpenStudy (anonymous):

excuse me i meant \[A'(t)=4\] not 40

OpenStudy (anonymous):

take the derivative of the formula \[A=\frac{1}{2}bh\] wrt time to get \[A'(t)=\frac{1}{2}(b'h + bh')\] via product rule

OpenStudy (anonymous):

we can solve for h' by \[2A'=b'h+bh'\] \[2A'-b'h=bh'\] \[h'=\frac{2A'-b'h}{b}\]

OpenStudy (anonymous):

got it, -2.2857 is my answer. thanks for your help :]

OpenStudy (anonymous):

yw.